Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Advanced Probability - Who Gets a Seat?

Expert replies
by beeparoo » Sat Jun 28, 2008 4:59 pm
Not only is it long, but it is complicated. Yet, despite knowing the answer, I have been stumped more than once. I welcome other unique approaches from various members.
++

Darcy, Gina, Ray and Susan will be the only participants at a meeting. There will be three soft chairs in the room where the meeting will be held, and one hard chair. No one can bring more chairs into the room. Darcy and Ray will arrive simultaneously, but Gina and Susan will arrive individually. The probability that Gina will arrive first is 1/3, and the probability that Susan will arrive first is 1/3. The probability that Gina will arrive last is 1/3, and the probability that Susan will arrive last is 1/3. Upon arriving at the meeting, each of the participants will select a soft chair, if one is available. If Darcy and Ray arrive and see only one unoccupied soft chair, they will flip a fair coin to determine who will sit in that chair. By what percent is the probability that Darcy will sit in a soft chair greater than the probability that Gina will sit in a soft chair?

A 50%
B 25%
C 16 2/3%
D 12 1/2%
E 0%

Source: Kaplan GMAT
Join the discussion
Source: — Problem Solving |

by wawatan » Sat Jun 28, 2008 7:36 pm
wow long problem.
Join the discussion

beeparoo wrote:Darcy, Gina, Ray and Susan will be the only participants at a meeting. There will be three soft chairs in the room where the meeting will be held, and one hard chair. No one can bring more chairs into the room. Darcy and Ray will arrive simultaneously, but Gina and Susan will arrive individually. The probability that Gina will arrive first is 1/3, and the probability that Susan will arrive first is 1/3. The probability that Gina will arrive last is 1/3, and the probability that Susan will arrive last is 1/3. Upon arriving at the meeting, each of the participants will select a soft chair, if one is available. If Darcy and Ray arrive and see only one unoccupied soft chair, they will flip a fair coin to determine who will sit in that chair. By what percent is the probability that Darcy will sit in a soft chair greater than the probability that Gina will sit in a soft chair?
The question is overly long, certainly longer than most real GMAT questions. The question is just trying to say that Gina, Susan and Darcy/Ray show up in random order- each has a 1/3 chance of showing up first, second, or last.

What's the probability that Gina will sit in a soft chair? If she shows up first or second, there will certainly be a soft chair available. If she shows up last, there won't. So 2/3.

What's the probability that Darcy will sit in a soft chair? If Darcy and Ray show up first or second, Darcy will get a soft chair: 2/3. If Darcy and Ray show up last, they flip a coin: 1/3 * 1/2 = 1/6. So the probability is 2/3 + 1/6 = 5/6 that Darcy gets a soft chair.

Finally, 5/6 is what percent greater than 2/3? (5/6 - 2/3)/(2/3) = 25%
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by beeparoo » Sun Jun 29, 2008 9:15 pm
Wow, Ian. Your confidance throughout your solution is unwavering. I hope to build that level of "but of course!" attitude on GMAT.

Anyway, I was unsure of my footing during the calculation of Darcy's probability and especially Gina's, thinking that I should treat two instances separately:

For calculation of Gina as second arrival, consider the possibility that
1) Susan enters first
or
2) Darcy/Gina enter first

Thoughtfully reviewing your response, I see, in retrospect, how they don't need to be quantified separately. But hindsight is 20/20 and clearly, I need to polish my probability skills.
Join the discussion